# A Comparative Analysis of Wealth Index Predictions in Africa between three Multi-Source Inference Modelsthanks: Please cite the ECMLPKDD’24 version of this paper

### v1

@@ 3 Results
Here, we compare the inferred poverty maps produced by M1, M2, and M3. We first present the predicted overall wealth distributions. Due to the lack of ground truth in all places of these poverty maps, we compare the predicted distributions with official GDP and Gini coefficient trends to understand how well these predictions align with established economic indicators. Given that each study employs different spatial resolutions, our comparison then focuses on locations that appear in all three studies. Note that these analyses are based on the predicted poverty maps rather than test sets with ground truth. We trained all variants of M3 models and use the best model in each country based on their lowest RMSE scores (see details in Section 2.2) to infer the poverty maps. The inferences from the other two models are publicly available: M1 [18], M2 [15]. For a comparison of the reported predictive power in terms of R2 for these studies on their respective test sets, please refer to Appendix Table A1.
@@ 3.1 Predicted wealth distribution
Table 2 summarizes the poverty map inferences for each country and model. First, we observe that M2 generally identifies more locations (populated places) than M1 and M3. Second, after re-scaling RWI scores for a fair comparison (see Equation 1), we find that M2 predicts the lowest mean wealth in most cases, except in South Africa. In contrast, the mean wealth predicted by M1 and M3 between each country is very similar. Third, in terms of standard deviation, M1 predicts a more homogeneous wealth distribution, characterized by lower values of SD and Gini coefficient. M3 follows, and M2 exhibits the highest SD and Gini coefficients, indicating greater wealth inequality in its predictions.
(a) High-resolution
(b) Districts
(c) IWI distribution
@@
Figure 2: IWI predictions in South Africa, ZAF (M3: CNNa+CB). All models were trained on the same ground-truth data (2016), thus, we expect similar results. Wealth predictions shown for M1, are linearly transformed into their IWI counterpart. Color bars display the average predicted IWI scores, centered at the mean of M3’s predictions for comparison. (a) Prediction at the original scale, 2.4 km grid-cells (M1), and OSM populated places (M2, M3). Both M2 and M3 include the Marion Island, omitted here for visualization purposes. (b) Aggregated mean wealth at the district level (administrative level 2). Compared to M3, M2 is consistently predicting higher wealth, while M1 predicts lower values. (c) Predicted wealth distribution (top: PDF, bottom: CDF). ZAF is the second country with the highest GDP in our study (see Figure 1), and this is reflected in the prediction by all models (e.g., $\mu>48$). Both M1 and M3 predict skewed distributions of wealth, while M2’s wealth follows an unusual multi-modal distribution. IWI poverty line ($35^{th}$ percentile, corresponding to poverty headcount ratios at \$1.25 a day [22]) is similar between M1 and M3, but differs for M2.
3.2 Expectation vs. Obtained Trends
@@
We compared M1 and M2 to M3, focusing on their expected performance based on changes in GDP per captia per country over the years of each model’s data. [^1] Since M3 is used as a baseline for comparison with both M1 and M2, comparing M1 and M2 directly is unnecessary because their relative performance can be inferred from their respective comparisons to M3. Results are shown in Table 3. In Uganda, where M2 and M3 used the same ground-truth, we expect similar wealth predictions. However, M2 inferred lower and more heterogeneous wealth than M3. In South Africa, despite all models using the same ground-truth data, we found differing results: M1 predicted slightly lower wealth values than M3, and M2 predicted slightly higher values. These patterns are visualized in Figure 2 (Section 0.A.2 for other countries). In Liberia, models M1, M2, and M3 used ground-truth data from $2013$, $2016$, and $2019/2022$, respectively (Table 1). During this period, GDP per capita rose from $717.6$ in $2013$ to $722.1$ in $2016$, dropped to $665.9$ in $2019$, and increased to $754.5$ in $2022$ (Figure 1). The normalized GDP difference between $2013$ and $2019$ ($0.04$) suggests that M3’s IWI scores should be slightly lower than M1’s, and that is what we obtained ($0.1$). However, even though the same trend was expected between M2 and M3, M2’s IWI scores were unexpectedly lower ($-0.23$). In terms of Gini coefficients, M3 generally predicted more heterogeneous IWI scores than M1 and more homogeneous scores than M2. In Uganda, M2 indicated greater inequality than M3 ($19.15$), which in turn showed more inequality than M1 ($-2.02$). The same pattern was observed in South Africa, although similar poverty skewness was expected across models. Due to data limitations, we were unable to track Gini coefficient trends for the other countries.
(a) M1 vs. M2
(b) M1 vs. M3
(c) M2 vs. M3
Figure 3: IWI predictions in overlapping locations in South Africa, ZAF (M3: CNNa+CB). Left: Correlation between predictions of two different models. Right: Poverty map using only overlapping locations. Colors indicate differences in predicted values per model pair: (a) M1 vs. M2, (b) M1 vs. M3, and (c) M2 vs. M3. M2 generally overestimates wealth compared to M1 (mostly red in (a)) and M3 (mostly blue in (c)). Moreover, the wealth distribution in these locations is equally skewed. In (b), M1’s predictions are more homogeneous than M3’s, despite achieving relatively low RMSE.
Table 4: Prediction differences in overlapping locations. Due to the differing spatial resolutions between the models (grid-cells vs. OSM populated places), we compare predictions only for overlapping locations. We spatially joined locations from M1 and M3, and M2 and M3, defining overlap as places in study $i$ that are also in study $j$ within a distance of less than $500m$. As expected, M3 contains more overlapping locations $N_{o}$ with M2 than with M1, since both M2 and M3 use OSM populated places. Although M2’s predictions are more positively correlated with M3’s, M3’s predictions are closer to M1’s (low RMSE) than to M2’s (high RMSE), consistent with the overall distributions (Table 2). Highlighted values accurately capture the expected trend’s sign (+/-)
Country M1 ($\hat{IWI}$) vs. M3 ($IWI$) M2 ($IWI$) vs. M3 ($IWI$)
@@
Diff.
Sierra Leone 1125 9.72 0.69 *** -0.16 -1.08 7264 6.89 0.87 *** 0.08 10.34
Liberia 1332 6.59 0.53 *** -0.11 -0.51 14609 11.06 0.74 *** 0.26 15.66
Uganda 2799 3.69 0.73 *** -0.02 0.06 13057 19.58 0.82 *** 0.26 20.47
Rwanda 95 4.63 0.25 * -0.01 0.27 1603 21.77 0.43 *** 0.20 21.19
South Africa 314 12.98 0.59 *** -0.04 -6.02 2866 22.78 0.45 *** -0.13 0.52
Gabon 43 6.47 0.62 *** -0.03 -3.29 598 25.21 0.73 *** 0.32 18.06
IWI Difference $=(\mu_{IWI_{M*}}-\mu_{IWI_{M3}})/(\mu_{IWI_{M*}}+\mu_{IWI_{M3}})$
Gini Difference $=(Gini_{M*}-Gini_{M3})$
3.3 Overlapping locations
The previous results compared the overall predictions across all models, providing a general understanding of their performance relative to country-level GDP and Gini coefficient trends. Now, we focus on comparing the predictions for locations that overlap in all three studies. It is important to note that the three studies differ in the spatial resolution of their poverty maps: M1 uses grid cells, while M2 and M3 use OSM populated places (see details in Section 2.1). Therefore, we perform a spatial join to analyze only these overlapping locations, which we expect to perfectly reflect the anticipated trends.
As before, we use M3 as a baseline model to compare the differences across all models, with results presented in Table 4. Unexpectedly, the observed trends are not consistently maintained across most countries and models. Notable exceptions include Uganda and Rwanda, where M3 infers higher mean wealth than M1, and Sierra Leone and Liberia, where M3 infers lower mean wealth than M2, aligning with the expected GDP trends over time (see Table 3). In the specific case of South Africa, where all models were trained using the same ground truth data, M3’s mean wealth in these overlapping locations is higher than that inferred by the other models. Nevertheless, M2’s heterogeneity is closer to that M3, while M1 produced more homogeneous wealth scores, see details in Figure 3.
@@ 4 Conclusions
To assess the alignment between the IWI score predictions by three models—M1 (Chi et al., 2021 [2]), M2 (Lee and Braithwaite, 2022 [16]), and M3 (Espín-Noboa et al., 2023 [8])—we examined the differences in their underlying ground-truth data and compared predicted wealth trends against expected changes in GDP per capita. Overall, the differences between M1 and M3 generally met expectations. Sierra Leone, Liberia, and Gabon showed higher mean wealth in M1 compared to M3, while South Africa showed lower values, and Uganda and Rwanda had similar ones. Small differences suggested stable wealth in Liberia, Uganda, Rwanda, and South Africa over time (M1 $\rightarrow$ M3). However, M2’s predictions did not align as expected, indicating lower poverty levels in several countries, except in South Africa, where M2’s IWI scores were slightly higher than M3’s. In general, the wealth distribution predicted by M3 was slightly more heterogeneous than that inferred by M1, but it was more homogeneous compared to M2. This trend was particularly expected and observed in Uganda between M1 and M3, and in South Africa when comparing M2 and M3.
When comparing the trends in the locations that appear in all three studies (i.e., overlaps), the expected trends were not consistently observed in most countries. Nevertheless, predictions by M1 are the closest to those by M3 compared to M2, particularly in Liberia, Uganda, Rwanda, South Africa and Gabon. Similar to the overall distributions, M2’s predictions for this set of places produce the most heterogeneous wealth distribution (i.e., highest inequality). Given the lack of official Gini coefficients for the years of interest for most countries, we cannot evaluate whether these trends in heterogeneity are expected. However, in South Africa, we anticipated minimal differences as all models use the same ground truth data. Only M2 and M3 reflect this expected trend.
These findings indicate that the predictions by Chi et al. and Espín-Noboa et al. align with general GDP trends over time. Differences in their predictions are expected since each study used data from different periods. For instance, Sierra Leone’s GDP decline from 2013 to 2016 and 2019 (Figure 1) is reflected in the corresponding drop in (re-scaled) RWI (2013) and IWI (2016, 2019) scores (Table 2). Discrepancies with the overall predictions by Lee and Braithwaite, as well as among the overlapping locations, underscore the necessity for policymakers and African stakeholders to further audit these models and predicted maps to ensure their validity before deployment. While evaluations during training are important, they are insufficient to capture the overall validity of the inferred wealth indicators, as they subjectively focus on the model’s performance on the test data. Accurate poverty maps are essential for effective policy interventions and play a crucial role in addressing the United Nations’ first sustainable development goal: eradicating poverty world-wide.
@@ 4.0.1 Acknowledgements
This work was funded by the SoBigData++ project (H2020-871042), Dataredux ANR project (ANR-19-CE46-0008), and the CHIST-ERA project SAI: FWFI 5205-N. The presented computational results have been achieved in part using the Vienna Scientifc Cluster (VSC). We thank the Data for Good team at Meta for insightful discussions during the first partner lightning talks, where early versions of this work were presented. We also appreciate the authors of the analyzed papers for their open data practices, which made this audit possible.
@@ Appendix 0.A Appendix
In this appendix, we present the comparison of the three models, M1, M2, and M3, across the remaining five African countries considered in this study: Sierra Leone, Liberia, Uganda, Rwanda, and Gabon. This analysis supplements our main findings by providing additional insights into the performance and predictions of each model, highlighting the differences in wealth estimates.
@@
0.A.1 Performance on test sets

### v3

@@ 3 Results
Here, we compare the inferred poverty maps produced by M1, M2, and M3. We first present the predicted overall wealth distributions. Due to the lack of ground truth in all places of these poverty maps, we compare the predicted distributions with official GDP and Gini coefficient trends to understand how well these predictions align with established economic indicators. Given that each study employs different spatial resolutions, our comparison focuses on locations that overlap within each pair of models (e.g., M1 vs. M2). Note that these analyses are based on the predicted poverty maps rather than test sets with ground truth. We trained all variants of M2 models and use the best model in each country based on their lowest RMSE scores (see details in Section 2.2) to infer the final poverty maps. The inferences of all models are publicly available: M1 [16], M2 [9], and M3 [21]. For a comparison of the reported predictive power in terms of R2 for these studies on their respective test sets, please refer to Section 0.A.2.
@@ 3.1 Predicted wealth distribution
Table 2 summarizes the poverty map inferences for each country and model. First, we observe that M1 generally identifies more locations (OSM populated places) compared to M2 and M3. Second, all models produce skewed distributions (indicating high wealth inequality), as reflected in their high Gini coefficients and significant skewness test. However, when comparing M1 with M2, we find that M1 generates more skewed distributions (higher Gini) and predicts lower mean wealth in most cases, except for South Africa.
(a) Poverty Map
(b) Wealth distribution
@@
Figure 2: IWI predictions in South Africa, ZAF (M2: CNNa+CB). Both models were trained on the same ground-truth data, thus, we expect similar results. (a) Poverty maps. Color bars display the average predicted IWI scores, centered at the mean of M2’s predictions for comparison. Top: Prediction at the original resolution, OSM populated places. Both M1 and M2 include the Marion Island, omitted here for visualization purposes. Bottom: Aggregated mean wealth at the district level (administrative level 2). Compared to M2, M1 predicts higher wealth. (b) Predicted wealth distribution (top: PDF, bottom: CDF). ZAF is the second country with the highest GDP in our study (see Figure 1), and this is reflected in the prediction by both models (e.g., $\mu>48$). M2 predicts a skewed, unimodal distribution of wealth, while M1 predicts an unusual multimodal distribution. IWI poverty line ($35^{th}$ percentile, corresponding to poverty headcount ratios at \$1.25 a day [27]) slightly differs between models.
3.2 Expected vs. Obtained Trends
@@
We compared M1 to M2, focusing on their expected performance based on changes in GDP per capita for each country over the years of each model’s data. [^1] We refer the reader to Section 0.A.3 for approximate comparisons with M3. Results are shown in Table 3. In Uganda, where both models used the same ground-truth, we expected similar wealth predictions. However, M1 inferred lower and more heterogeneous wealth compared to M2. A similar expectation applied to South Africa, but while the results were comparable on average, M1 predicted wealthier places and more heterogeneous. In Liberia, M1 used $2016$ ground-truth data, while M2 used data from $2019$ and $2022$ (Table 1). During this period, GDP per capita dropped from $722.1$ in $2016$ to $665.9$ in $2019$, then rose to $754.5$ in $2022$ (Figure 1). Averaging GDP between $2019$ and $2022$ (years used to train M2), we expected a slight decrease from $722.1$ to $710.2$. This is reflected in the normalized GDP difference ($0.01$), suggesting that M2’s IWI scores should be slightly lower than M1’s. However, M2 unexpectedly predicted higher values ($-0.23$). In terms of Gini coefficients, M1 generally predicted more heterogeneous wealth distributions than M2. Due to data limitations, we were unable to track Gini coefficient trends for the other countries. These patterns are visualized in Figure 2 for South Africa, and Section 0.A.1 for other countries.
Table 4: Prediction differences in overlapping places. Predictions are compared only for locations common to both models, defined as those within $500~{}m$ of each other. Duplicates were removed by retaining the closest pairs. As expected, M2 contains many locations from M1, as both use OSM populated places, though M1 includes more (Table 2). Predictions are most similar in SLE (low RMSE, high Pearson correlation), with larger differences observed in other countries. Highlighted values show the correct trend (+/-).
Country’s Code M1 vs. M2
@@
SLE 6227 47.75 63.02 6.63 0.80 *** 0.00 -0.10 NA 8.89
LBR 13786 83.43 88.39 11.00 0.68 *** 0.01 -0.27 NA 14.46
UGA 10981 39.96 39.51 19.96 0.78 *** 0.00 -0.28 0 19.31
RWA 653 2.80 56.78 22.18 0.35 *** -0.05 -0.22 NA 21.04
ZAF 2686 11.84 88.27 22.61 0.46 *** 0.00 0.13 0 0.91
GAB 581 12.02 51.19 25.13 0.73 *** NA -0.32 NA 18.21
Figure 3: IWI predictions at overlapping places in South Africa, ZAF (M2: CNNa+CB). Right: IWI differences between the predictions of M1 and M2 for each overlapping place. Left: Correlation between these predictions. Colors indicate the direction of the differences: M1 generally overestimates wealth compared to M2 (mostly blue). Moreover, the wealth distribution in these locations is equally skewed (similar Gini coefficient).
3.3 Overlapping places
The previous results compared overall predictions across models, offering a broad understanding of their performance relative to country-level GDP per capita and Gini coefficient trends. Now, we focus on comparing the predictions for locations that overlap in both M1 and M2. We perform a spatial join to analyze only these overlapping locations, which we expect to perfectly reflect the anticipated trends as both models rely on OSM populated places (see Section 2.1).
As before, we compare the expected GDP per capita differences with the observed variations in predicted mean wealth between the two models, see Table 4. When comparing these findings with the overall predictions (Table 3), we observe consistent trends in both wealth predictions and Gini coefficients. In all countries, M2 predicts higher average wealth than M1, with the exception of ZAF. Additionally, M1 consistently produces more skewed distributions, indicating higher wealth inequality. However, the magnitude of these differences is generally smaller in overlapping locations compared to all locations (Table 3). This suggests that while the trends in overlapping locations do not perfectly match expectations, the values are closer to what was anticipated. For instance, in UGA and ZAF, where identical predictions were expected, the results align more closely with the anticipated outcome. In RWA, as predicted, M2 identifies wealthier places, but with a slight over-estimation. The correlation between these predicted wealth scores for ZAF is shown in Figure 3 (Section 0.A.1 for other countries).
@@ 4 Conclusions
This study aimed to evaluate the alignment between predicted wealth trends from three models—M1 (Lee and Braithwaite, 2022 [17]), M2 (Espín-Noboa et al., 2023 [8]), and M3 (Chi et al., 2022 [2])—using IWI and RWI scores as proxies for economic well-being. The comparison of these models centered on understanding their differences in underlying ground-truth data and their ability to predict wealth in relation to expected changes in GDP per capita and wealth inequality (Gini coefficient).
Our primary focus was on comparing IWI predictions from M1 and M2, as they directly provide comparable outputs. While M3 predicts RWI scores, the conversion to IWI for proper comparison remains complex and non-trivial. Although some prior studies [25, 28] have proposed methods for these conversions, they come with inherent limitations. For the sake of simplicity and based on the high correlation between RWI and IWI scores (Pearson’s correlation: $0.93-0.99$ [27]), we included M3 in our comparison (see Section 0.A.3), but the core analysis remained on M1 and M2.
Our findings indicate unexpected differences between the predictions of M1 and M2. M2 generally predicts wealthier regions, while M1 over-represents poverty. Based on GDP per capita trends during the training years, the models should produce similar wealth predictions, except in Rwanda, where the trend direction is captured but exaggerated. In Uganda and South Africa, where each model used the same years of ground-truth data, we expected similar wealth scores. However, M1 predicted lower wealth in Uganda and higher wealth in South Africa, contradicting our hypothesis of consistent distributions between the models.
In places where both models overlap, observed trends from the predictions also deviated from expected wealth trajectories (Table 4). However, these discrepancies were smaller compared to the ones observed in the overall predictions (Table 3). In the overlapping places, M1 remained more heterogeneous than M2, suggesting greater inequality. The absence of official Gini coefficients for the relevant years prevented us from verifying whether the observed wealth inequality trends in Sierra Leone, Liberia, Rwanda, and Gabon reflect reality.
The observed discrepancies between model predictions highlight the critical need for further scrutiny before using these inferred poverty maps for policy-making. While model evaluations during training provide insight into test data performance, they do not guarantee the broader accuracy of the inferred wealth indicators. Accurate and reliable poverty maps are fundamental to informing effective policy interventions, particularly in line with the United Nations’ first Sustainable Development Goal: eradicating global poverty.
Moving forward, more comprehensive audits of these models and their underlying assumptions are essential. Collaboration between AI practitioners, policymakers, and local stakeholders in African countries is crucial to ensure that poverty maps accurately reflect local realities. Future research should prioritize refining methodologies for converting RWI to IWI scores to enable more nuanced comparisons across wealth indices. Alternatively, a replication study using M3’s architecture to predict IWI scores could offer more precise cross-model cross-country and cross-year comparisons. Governments, policymakers, and NGOs must prioritize providing up-to-date global statistics to enhance the evaluation of wealth inequality trends generated by machine learning and AI technologies, thereby supporting progress toward broader Sustainable Development Goals (SDGs).
@@ 4.0.1 Acknowledgements
This work was funded by the SoBigData++ project (H2020-871042), Dataredux ANR project (ANR-19-CE46-0008), and the CHIST-ERA project SAI: FWFI 5205-N. The presented computational results have been achieved in part using the Vienna Scientifc Cluster (VSC). This work was partially funded by the Vienna Science and Technology Fund WWTF under project No. ICT20-079. We thank the Data for Good team at Meta for insightful discussions during the first partner lightning talks, where early versions of this work were presented. We also appreciate the authors of the analyzed papers for their open data practices, which made this audit possible.
@@ Appendix 0.A Appendix
This appendix provides additional material comparing models M1 and M2 across five African countries: Sierra Leone, Liberia, Uganda, Rwanda, and Gabon. Due to the presence of negative RWI values in M3 and the complex conversion to IWI scores [25], we compare the distribution shapes across all models and assess their similarity, along with the reported validation scores. This analysis complements our main findings by offering further insights into the models’ performance and wealth estimates.
@@
0.A.1 M1 vs. M2 on remaining countries
In this section, we present the plots comparing the predictions of models M1 and M2 across the remaining five African countries, complementing the results presented in the main text.
0.A.1.1 Sierra Leone, SLE
(a) Poverty Map
(b) Wealth distribution
Figure A1: IWI predictions in SLE (M2: CB). (a) Poverty maps with colors representing the average predicted IWI scores. Top: OSM populated places. Bottom: District level. M1 predicts lower wealth than M2. (b) Predicted wealth distribution. IWI poverty line ($35^{th}$ percentile, corresponding to poverty headcount ratios at \$1.25 a day [27]) slightly differs between models.
Figure A2: IWI predictions in overlapping places in SLE (M2: CB). Right: Prediction differences between M1 and M2. Left: Correlation between these predictions. Compared to M2, M1 underestimates wealth (mostly red) and predicts more inequality (higher Gini coefficient).
0.A.1.2 Liberia, LBR
(a) Poverty Map
(b) Wealth distribution
Figure A3: IWI predictions in LBR (M2: CB). (a) Poverty maps with colors representing the average predicted IWI scores. Top: OSM populated places. Bottom: District level. M1 predicts lower wealth than M2. (b) Predicted wealth distribution. IWI poverty line ($35^{th}$ percentile, corresponding to poverty headcount ratios at \$1.25 a day [27]) slightly differs between models.
Figure A4: IWI predictions in overlapping places in LBR (M2: CB). Right: Prediction differences between M1 and M2. Left: Correlation between these predictions. Compared to M2, M1 underestimates wealth (mostly red) and predicts more inequality (higher Gini coefficient).
0.A.1.3 Uganda, UGA
(a) Poverty Map
(b) Wealth distribution
Figure A5: IWI predictions in UGA (M2: CBw). (a) Poverty maps with colors representing the average predicted IWI scores. Top: OSM populated places. Bottom: District level. M1 predicts lower wealth than M2. (b) Predicted wealth distribution. IWI poverty line ($35^{th}$ percentile, corresponding to poverty headcount ratios at \$1.25 a day [27]) differs between models.
Figure A6: IWI predictions in overlapping places in UGA (M2: CBw). Right: Prediction differences between M1 and M2. Left: Correlation between these predictions. Compared to M2, M1 underestimates wealth (mostly red) and predicts more inequality (higher Gini coefficient).
0.A.1.4 Rwanda, RWA
(a) Poverty Map
(b) Wealth distribution
Figure A7: IWI predictions in RWA (M2: CBw). (a) Poverty maps with colors representing the average predicted IWI scores. Top: OSM populated places. Bottom: District level. M1 predicts lower wealth than M2. (b) Predicted wealth distribution. IWI poverty line ($35^{th}$ percentile, corresponding to poverty headcount ratios at \$1.25 a day [27]) differs between models.
Figure A8: IWI predictions in overlapping places in RWA (M2: CBw). Right: Prediction differences between M1 and M2. Left: Correlation between these predictions. Compared to M2, M1 underestimates wealth (mostly red) and predicts more inequality (higher Gini coefficient).
0.A.1.5 Gabon, GAB
(a) Poverty Map
(b) Wealth distribution
Figure A9: IWI predictions in GAB (M2: CNNa+CB). (a) Poverty maps with colors representing the average predicted IWI scores. Top: OSM populated places. Bottom: District level. M1 predicts lower wealth than M2. (b) Predicted wealth distribution. IWI poverty line ($35^{th}$ percentile, corresponding to poverty headcount ratios at \$1.25 a day [27]) differs between models.
Figure A10: IWI predictions in overlapping places in GAB (M2: CNNa+CB). Right: Prediction differences between M1 and M2. Left: Correlation between these predictions. Compared to M2, M1 underestimates wealth (mostly red) and predicts more inequality (higher Gini coefficient).
0.A.2 Performance on test sets
In Table A1, we present the reported evaluation scores (R2) on the test set for each model across different countries, provided solely for background context. It is important to note that these scores are not directly comparable due to several factors. First, each model was trained on data from different years and places, leading to variations in the socioeconomic conditions captured. Even in cases like UGA and ZAF, where M1 and M2 used data from the same year, their training, validation, and test sets were not identical. Second, while M1 and M2 predict IWI (International Wealth Index) scores, M3 predicts RWI (Relative Wealth Index) scores, further complicating a direct comparison. Therefore, the R2 values in Table A1 should be seen as individual performance metrics rather than for cross-model comparison.